OpenStudy (anonymous):

A right circular cylinder has a base whose area is 616 square feet. If the height of the cylinder is 9 feet, what is its lateral surface area? use pi = 22/7. 1408 ft2 396 ft2 792 ft2 5544 ft2

OpenStudy (anonymous):

I tried drawing it out and assumed 616 was the diameter, and I used 308 for the radius in the formula for the lateral area of a cylinder, but I got 17416 as an asnwer :/

OpenStudy (ganpat):

Lateral surface area = 2 * pi * r * h = 2 * 22/7 * r * 9 -----> 1 2. Area at base = pi * r2 = 22/7 * r2 = 616 so r = 14 ft.... substitute in 1.. So Lateral area = 2 * 22/7 * 14 * 9 = 792 sq ft... option 3...